OpenStudy (anonymous):

how do you find the volume of the solid. a sphere with the diameter of 3 and 1/2

OpenStudy (anonymous):

Volume of a sphere = \[4/3 \pi r^3\] \[Diameter = 2 radius\] So\[ r= d/2\]